# Motion-R1: Chain-of-Thought Reasoning and Reinforcement Learning for Human Motion Generation

### v3

@@ Abstract
Recent advances in large language models, especially in natural language understanding and reasoning, have opened new possibilities for text-to-motion generation. Although existing approaches have made notable progress in semantic alignment and motion synthesis, they often rely on end-to-end mapping strategies that fail to capture deep linguistic structures and logical reasoning. Consequently, generated motions tend to lack controllability, consistency, and diversity. To address these limitations, we propose Motion-R1, a unified motion-language modeling framework that integrates a Chain-of-Thought mechanism. By explicitly decomposing complex textual instructions into logically structured action paths, Motion-R1 provides high-level semantic guidance for motion generation, significantly enhancing the model’s ability to interpret and execute multi-step, long-horizon, and compositionally rich commands. To train our model, we adopt Group Relative Policy Optimization, a reinforcement learning algorithm designed for large models, which leverages motion quality feedback to optimize reasoning chains and motion synthesis jointly. Extensive experiments across multiple benchmark datasets demonstrate that Motion-R1 achieves competitive or superior performance compared to state-of-the-art methods, particularly in scenarios requiring nuanced semantic understanding and long-term temporal coherence. The code, model and data will be publicly available.
@@ 1 Introduction
Text-to-Motion (T2M) generation guoBackMLPSimple2022a ; tevet2022human ; Guo_2022_CVPR has emerged as a fundamental task in human-machine interaction, enabling the synthesis of realistic human motions from natural language descriptions. Driven by rapid advancements in large language models (LLMs) achiam2023gpt ; li2023blip ; qi2024cocogesture , recent T2M approaches  jiangMotionGPTHumanMotion2023 ; wangMotionGPT2GeneralPurposeMotionLanguage2024 ; wang2024motionagent have made significant strides in generating high-fidelity motions that align with complex textual instructions. Despite this progress, current approaches predominantly rely on end-to-end supervised learning ahnText2ActionGenerativeAdversarial2017 ; hongAvatarCLIPZeroShotTextDriven2022 ; ahujaLanguage2PoseNaturalLanguage2019 , mapping textual inputs directly onto motion sequences. While straightforward, such methods often fail to capture deeper linguistic structures and complex logical relationships embedded in natural language. For example, a high-level instruction such as "a person makes a cup of coffee" implicitly involves a sequence of semantically connected sub-actions, such as reaching, grasping, pouring, stirring, and placing, which require temporal ordering and causal reasoning. End-to-end models often struggle to interpret and execute such compositional commands accurately, resulting in incoherent or oversimplified motions.
Inspired by the recent success of Chain-of-Thought (CoT) weicot2022 prompting in enhancing reasoning capabilities of LLMs, we hypothesize that explicitly modeling intermediate reasoning steps can similarly benefit motion generation tasks. By decomposing high-level instructions into structured action plans, models can better capture temporal dependencies, hierarchical goals, and fine-grained semantic nuances embedded in language. Moreover, reinforcement learning (RL) kaelbling1996reinforcement , especially Group Relative Policy Optimization (GRPO) shao2024deepseekmathpushinglimitsmathematical , offers a natural framework for refining such reasoning and generation processes by optimizing them for downstream motion quality. As shown in Figure 1, combining CoT-style semantic decomposition with RL-based optimization can improve the model’s generalization to out-of-distribution inputs, enhance the interpretability of action sequences, and facilitate more precise control over complex motion synthesis.
To this end, we propose Motion-R1, a unified T2M framework that integrates a CoT mechanism. By combining structured reasoning with reward-driven optimization, Motion-R1 significantly enhances the model’s generalization ability, capability to explore diverse motion styles, and interpretability of generated motions.
Specifically, we propose the MotionCoT Data Engine, an automated pipeline that leverages LLMs to generate high-quality Chain-of-Thought (CoT) annotations for motion descriptions. By distilling the semantic planning abilities of LLMs into structured action guidance, it helps the model capture the underlying logic of human actions. This enriched supervision improves generalization in motion synthesis. Compared to manual annotation, our approach produces more diverse and consistent CoT data at lower cost, serving as both an effective training signal and an interpretable bridge between text and motion.
We extend GRPO to the motion generation domain, offering a lightweight and scalable alternative to conventional RL. GRPO requires no value networks (e.g., PPO) or human preference data (e.g., DPO rafailov2023direct , RLHF), and instead uses group-wise ranking to learn from quality differences in generated samples—enhancing both generalization and diversity. To improve semantic alignment, we design a multi-aspect reward tailored to motion synthesis and adopt supervised fine-tuning (SFT) as a cold-start strategy for stable RL training.
The contributions of our work are summarized as follows:
$\bullet$ We introduce Motion-R1, a reasoning-guided and reinforcement-optimized framework for text-to-motion generation. It enhances the model’s ability to interpret abstract and complex textual descriptions through CoT planning and improves generalization to out-of-distribution instructions.
$\bullet$ We develop MotionCoT Data Engine, an automated CoT annotation pipeline that leverages LLMs to generate structured motion planning paths. This enables effective distillation of LLMs’ reasoning capabilities and reduces reliance on costly manual labeling.
$\bullet$ We extend the GRPO algorithm to the motion generation domain, offering a simple yet effective reinforcement learning paradigm that requires no value networks or preference data. Additionally, we propose multi-aspect reward functions for motion synthesis to strengthen semantic grounding and training stability.
$\bullet$ Extensive experiments demonstrate that Motion-R1 achieves state-of-the-art performance on multiple benchmarks. Compared with MotionAgent wang2024motionagent , our method exhibits superior semantic understanding and motion synthesis capabilities, particularly for complex and out-of-distribution textual descriptions. Specifically, Motion-R1 outperforms MotionAgent by achieving significant improvements of $\mathbf{1.19}\%$ in motion diversity and $\mathbf{2.12}\%$ in R-precision.
@@ 2 Related Work
2.1 Multimodal Large Language Models
@@
Recently, multimodal large language models (MLLMs) wu2023multimodal ; koh2023generating ; tang2025video have shown impressive performance in various tasks, including text generation, reasoning, and even multimodal tasks. For example, models like GPT-4 achiam2023gpt and BLIP-2 li2023blip have demonstrated the ability to understand and generate text and images. While recent MLLMs have significantly improved their perception and generation capabilities across diverse modalities, many complex real-world tasks demand not only understanding but also reasoning banerjeeWeaklySupervisedRelative2021 ; xiongTEILPTimePrediction2024 . Consequently, enhancing the reasoning abilities of MLLMs has become an important research direction.
Building on the success of CoT  weicot2022 prompting in NLP, researchers have explored extending CoT-style reasoning to multimodal settings. For instance, models like LLaVA-CoT xuLLaVACoTLetVision2025 introduce a novel vision-language model capable of performing autonomous, structured, and multistage reasoning by decomposing complex questions into four stages: summary, caption, reasoning, and conclusion. While structured reasoning improves the interpretability and controllability of vision-language models, recent studies tan2025reason ; pan2025medvlm ; liu2025seg further suggest that reasoning capabilities can be learned and optimized via feedback-driven training. DeepSeek-R1 deepSeekR1IncentivizingReasoning2025 introduces GRPO, a reinforcement learning algorithm that trains LLMs to produce more accurate and coherent responses, leveraging reward signals from downstream tasks. Although this approach has shown promise in improving the reasoning capabilities of MLLMs huangVisionR1IncentivizingReasoning2025 , its application in the context of 3D human motion generation remains unexplored. This idea inspires our approach, which extends GRPO to the domain of motion generation, using action quality as feedback to refine both the thought process and the generated motion sequence.
2.2 Text-guided 3D Human Motion Generation
@@
Text-guided 3D human motion generation has become a key research focus in recent years. Early methods ahnText2ActionGenerativeAdversarial2017 ; ghoshSynthesisCompositionalAnimations2023 ; qi2024weakly primarily relied on generative adversarial networks goodfellow2014generative ; wang2024hallo3d to synthesize human motion from text descriptions. These methods often struggled with issues such as limited diversity.
@@
To address these challenges, diffusion models ho2020denoising ; li2024era3d ; li2025pshuman ; wang2025humandreamergeneratingcontrollablehumanmotion ; wang2025humandreamerxphotorealisticsingleimagehuman have gained popularity in this field. Notable works zhouUDEUnifiedDriving2022 ; kimFLAMEFreeformLanguagebased2023 ; dabralMoFusionFrameworkDenoisingDiffusionbased2023 include MDM (tevet2022human, ), which first applies diffusion modeling to motion synthesis; MotionDiffuse (zhang2023motiondiffuse, ) improves temporal consistency; MotionChain (jiang2023motionchain, ) explores fine-grained structure and staged modelling; and MoMask (guo2023momask, ) employs masking strategies for improved training efficiency. However, these methods often rely on complex architectures and extensive training data and are limited by their motion initialization.
@@
In parallel, VQ-VAE-based methods Guo_2022_CVPR ; hongAvatarCLIPZeroShotTextDriven2022 ; petrovichTEMOSGeneratingDiverse2022 ; guoTM2TStochasticTokenized2022 ; athanasiouTEACHTemporalAction2022 discretize motion representations to facilitate efficient sequence modeling and better alignment with language, providing a more interpretable latent space. Building on this, LLMs have been integrated into the generation process to enhance the quality and control of generated motions. For instance, T2M-GPT (zhangT2MGPTGeneratingHuman2023, ) combines VQ-VAE with GPT in a two-stage framework for high-quality generation. Subsequent works such as MotionDiffuse zhangMotionDiffuseTextDrivenHuman2022 , MotionCLIP (tevet2022motionclip, ), and MotionGPT (jiangMotionGPTHumanMotion2023, ) integrate diffusion modeling, contrastive learning, and Transformer architectures to enhance fidelity and semantic consistency further. MotionAgent (wang2024motionagent, ) further incorporates semantic planning for robust generalization in complex task scenarios. Meanwhile, with the development of reinforcement learning, MotionRL liuMotionRLAlignTexttoMotion2024 leverages PPO schulman2017proximal to improve the generation based on human preferences prior knowledge.
However, despite these advances, current T2M methods largely adopt an end-to-end mapping mode, lacking explicit reasoning capabilities over complex linguistic instructions. Our approach aims to address this gap by introducing a CoT-guided and reward-optimized framework for 3D human motion generation, enhancing the interpretability and diversity of the generated motions.
@@ 3 Method
Figure 2: Motion-R1 pipeline. The framework consists of two stages: (1) MotionCoT Data Engine uses DeepSeek-R1 to generate CoT-style motion planning traces in <think>, <output>, and <Motion> format to fine-tune an LLM; (2) GRPO-based training ranks grouped outputs by format, motion, and semantic rewards to optimize the LLM via reinforcement learning.
@@ 3.1 Motion-R1 Framework
Motion-R1 comprises two core components: a pre-trained motion tokenizer and an LLM equipped with action-oriented reasoning capabilities. The motion tokenizer discretizes continuous motion sequences into motion tokens and reconstructs them back into smooth, coherent trajectories. On the other hand, the LLM is designed to perform structured reasoning over natural language instructions, enabling it to decompose complex action descriptions into finer-grained, logically ordered sub-actions — what we refer to as a reasoning-based motion breakdown. Based on this structured interpretation, the model generates high-quality motion token sequences that faithfully reflect the intended behavior.
As shown in Figure 2, our framework consists of two training stages. In the first stage, we construct a motion-language dataset with CoT supervision weicot2022 by prompting DeepSeek-R1 deepSeekR1IncentivizingReasoning2025 . This dataset, referred to as the MotionCoT Data Engine, enables cold-start supervised tuning of the LLM to produce structured reasoning process in the <think>, <output>, and <Motion> format.
In the second stage, we refine the LLM via reinforcement learning using GRPO shao2024deepseekmathpushinglimitsmathematical . GRPO operates by sampling $G$ generations per prompt and ranking them within a group. For each sampled output, a set of task-specific rewards $R_{n}$ is computed, including format correctness, motion similarity, and semantic alignment. Based on these rankings, relative advantages $A_{n}$ are estimated to update the policy without requiring an explicit value function. These feedback signals guide the LLM to generate more accurate and realistic motion sequences conditioned on textual instructions.
@@ 3.2 Motion Tokenizer
To integrate motion data, which differs significantly from natural language in structure and modality, into the LLM framework, we adopt a VQ-VAE architecture as our motion tokenizer. This approach has been widely adopted in  zhangT2MGPTGeneratingHuman2023 ; jiangMotionGPTHumanMotion2023 ; guo2023momask ; wang2024motionagent and proven effective for 3D human motion data modeling.
The motion tokenizer comprises an encoder $E$ and a decoder $D$. Given an input motion sequence $\mathbf{m}_{1:T}\in\mathbb{R}^{T\times D}$, where $T$ denotes the number of frames and $D$ is the number of dimensions of each frame, the encoder $E$ maps the sequence to a latent representation $\mathbf{z}_{1:(T/l)}\in\mathbb{R}^{(T/l)\times d}$, where $l$ is the temporal downsampling rate and $d$ is the number of latent dimensions. Each latent vector $\mathbf{z_{i}}$ is then quantized using a learnable codebook $\mathbf{C}=\{\mathbf{c}_{n}\}_{n=1}^{N}$, where $N$ is the codebook size and $\mathbf{c}_{n}\in\mathbb{R}^{d}$ represents a discrete motion token. The quantization selects the nearest code vector to each embedding:
@@
The original motion sequence is reconstructed as $\hat{\mathbf{m}}_{1:T}=D(\hat{\mathbf{z}}_{1:(T/l)})$. Following zhangT2MGPTGeneratingHuman2023 , we train the VQ-VAE model using a composite objective that includes a reconstruction loss $L_{\text{reconstruct}}$, a codebook commitment loss $L_{\text{commit}}$, and an embedding loss $L_{\text{embed}}$:
@@
Here, $L_{\text{reconstruct}}$ includes a smoothed L1 loss with velocity regularization to improve generation quality. To ensure stable and efficient training, we follow  zhangT2MGPTGeneratingHuman2023 to incorporate exponential moving average (EMA) updates for the codebook along with codebook reset strategies.
Figure 3: MotionCoT Data Example. Given the prompt “A person does Tai chi.” the LLM generates a step-by-step CoT reasoning trace (<think>) and a structured action plan <output>, covering stance, arm movement, weight transfer, and hand positioning.
3.3 MotionCoT Data Engine
We present MotionCoT Data Engine, an automated framework for constructing CoT weicot2022 supervision tailored to the text-to-motion generation task. This engine leverages the general reasoning and semantic planning capabilities of LLM to transform free-form motion descriptions into structured, high-quality CoT action plans, which serve as intermediate guidance between language and motion.
The data engine begins by augmenting existing motion-language datasets through prompt-based querying. We design a comprehensive prompt that instructs the LLM to reason through the detailed logic and temporal structure. This prompt includes clear task instructions, output format constraints, and several in-context examples to guide the model in producing structured, step-by-step CoT reasoning. The detailed prompt is provided in the supplementary material.
The CoT outputs generated by this process are then passed through an automated quality control pipeline. Specifically, we employ DeepSeek-R1 deepSeekR1IncentivizingReasoning2025 to evaluate each generated CoT for relevance, conciseness, and logical coherence. Samples exhibiting redundancy, overthinking, or excessive verbosity are filtered out and regenerated iteratively until they meet our quality standards. As shown in Figure 3, we give an example of the CoT generation process. The LLM is prompted to generate a CoT reasoning process for the task of “A person does Tai chi”. The model first identifies the main action, “Tai chi”, and then breaks it down into sub-actions such as “stand up”, “arm movements”, “weight displacement”, and “hand positioning”. Each sub-action is further elaborated with specific details, such as the direction of movement and the body parts involved.
Once validated, each high-quality CoT trace is paired with its original motion description and corresponding action sequence, forming triplets of the form (description, CoT, motion). These triplets constitute a refined dataset that provides explicit, interpretable reasoning supervision for downstream training.
By distilling structured action plans from LLMs, the MotionCoT Data Engine enables the model to learn fine-grained and context-aware semantic decomposition from natural language. This significantly enhances the model’s controllability, interpretability, and generalization capability in complex motion generation tasks while dramatically reducing the cost of manual annotation.

### v4

# Motion-R1: Enhancing Motion Generation with Decomposed Chain-of-Thought and RL Binding

@@ Abstract
Text-to-Motion generation has become a fundamental task in human-machine interaction, enabling the synthesis of realistic human motions from natural language descriptions. Although recent advances in large language models and reinforcement learning have contributed to high-quality motion generation, two major challenges remain. Existing approaches often fail to capture the temporal and causal complexities inherent in natural language, leading to oversimplified or incoherent motions. Additionally, RL-based methods are frequently overly complex, hindering their scalability and adaptability across various motion generation tasks. To address these challenges, we propose Motion-R1, a novel framework that combines decomposed Chain-of-Thought reasoning with reinforcement learning to enhance both the quality and interpretability of generated motions. Specifically, we introduce the Decomposed CoT Data Engine, which leverages an automated pipeline to synthesize high-quality reasoning data, allowing the model to better capture the temporal dependencies and causal relationships of human motion. We also propose RL Binding, a reinforcement learning strategy that incorporates multi-modal text-motion alignment into the RL reward function, guiding the model to produce motions that are both semantically accurate and motionally realistic. Extensive experiments across benchmark datasets demonstrate that Motion-R1 achieves state-of-the-art performance, with a 3.5% improvement in MM-Dist on HumanML3D and improvements in R-Precision and FID on KIT-ML and BABEL, surpassing existing methods across key metrics and highlighting its superior capability in handling complex motion generation tasks. Project page: https://motion-r1.github.io/.
@@ 1 Introduction
Text-to-Motion (T2M) generation Guo et al. (2022c); Tevet et al. (2023); Guo et al. (2022a) has emerged as a fundamental task in human-machine interaction, enabling the synthesis of realistic human motions from natural language descriptions. Driven by rapid advancements in large language models (LLMs) Achiam et al. (2023); Li et al. (2023), recent T2M approaches  Jiang et al. (2023); Wang et al. (2024); Wu et al. (2024) have made significant strides in generating high-fidelity motions that align with complex textual instructions. Reinforcement learning (RL) Kaelbling et al. (1996) provides a promising approach to enhance motion generation by optimizing it for motion quality. Recent works Liu et al. (2024); Haoru Wang et al. (2025) have successfully integrated RL into motion generation, improving both text adherence and motion quality by aligning with human perceptual preferences. However, despite significant advancements in motion generation, current methods still face two major challenges.
(1) Existing approaches predominantly rely on end-to-end supervised learning Ahn et al. (2017); Hong et al. (2022); Ahuja & Morency (2019), directly mapping textual inputs to motion sequences. While this approach is simple, it fails to capture the deeper temporal and causal relationships inherent in natural language. For instance, a high-level task like ”making a cup of coffee” involves a series of connected sub-actions (e.g., reaching, grasping, pouring, stirring, placing), which require careful temporal ordering and causal reasoning. However, these methods often struggle to effectively decompose such tasks, resulting in oversimplified or incoherent motion generation.
(2) RL-based methods, such as MotionRL Liu et al. (2024) and MotionCritic Haoru Wang et al. (2025), demonstrate improvements in motion quality but are often overly complex and over-engineered. These designs, though effective, limit their adaptability to real-world applications. The intricate nature of these RL models makes them difficult to scale and deploy across a wide range of motion generation tasks, especially when simplicity and efficiency are needed.
Figure 1: Comparison of traditional approaches and our Motion-R1 framework. (a) Traditional end-to-end models exhibit poor generalization on out-of-distribution motions. (b) Our Decomposed CoT Data Engine enables strong generalization by structuring high-level instructions into intermediate reasoning steps. (c) Existing RL-based methods rely on expensive human annotations to train preference models for reward signals. (d) Our RL Binding mechanism achieves efficient multi-modal alignment without additional annotation cost.
Our motivation is to advance motion generation by introducing a novel framework, Motion-R1 that effectively addresses the key challenges in the field, as shown in Figure 1.
To tackle the first challenge, we propose a Decomposed CoT Data Engine which synthesizes high-quality, step-by-step reasoning data. Inspired by the recent success of Chain-of-Thought (CoT) Wei et al. (2022) prompting in enhancing reasoning capabilities of LLMs, we hypothesize that explicitly modeling intermediate reasoning steps can similarly benefit motion generation tasks. By decomposing high-level instructions into structured action plans, our engine leverages an automated CoT annotation pipeline that employs LLMs to generate structured motion planning paths, enabling models to better capture temporal dependencies, causal relationships, and fine-grained semantic nuances embedded in language. This approach not only enhances the realism and coherence of the generated motions but also reduces reliance on expensive manual annotation, making the process more efficient and scalable.
To address the second challenge, we introduce RL Binding, a novel reinforcement learning strategy that enhances motion generation by incorporating a multi-modal alignment mechanism within the RL framework. RL Binding simultaneously aligns the generated motion sequences with both the ground-truth motions and the corresponding textual descriptions. This dual alignment process ensures that the generated motions not only stay temporally consistent with real-world actions but also faithfully capture the semantic meaning conveyed in the textual instructions. By embedding these alignments directly into the RL reward function (specifically, the motion similarity reward, the semantic similarity reward, and the format reward), RL Binding guides the model to produce motions that are both semantically accurate and motionally realistic. This streamlined design simplifies the optimization process, improving both the quality and interpretability of the generated motions, while offering a highly adaptable and efficient solution for real-world applications.
Moreover, To demonstrate the effectiveness of our approach, we conduct extensive experiments on multiple benchmark datasets. Our method achieves a 3.5% improvement in MM-Dist on HumanML3D Guo et al. (2022a), while also setting new state-of-the-art records in R-Precision and FID on both KIT-ML Plappert et al. (2016) and BABEL Punnakkal et al. (2021). These results highlight the superiority of our method in both motion quality and task performance, showcasing its strong potential for real-world applications.
In general, our work’s contributions can be summarized in the following three folds:
$\bullet$ Decomposed CoT Data Engine: We introduce a novel framework for synthesizing high-quality, step-by-step reasoning data. By leveraging an automated CoT annotation pipeline powered by LLMs, our approach effectively captures the temporal dependencies and causal relationships inherent in motion generation. This method significantly improves the quality of generated motions while reducing the reliance on costly manual annotation, making the process more efficient and scalable.
$\bullet$ RL Binding Strategy: We propose RL Binding, a novel reinforcement learning strategy that integrates multi-modal text-motion alignment into the reward function via the motion similarity reward, the semantic similarity reward, and the format reward, effectively guiding the model to generate motions that better adhere to textual instructions while maintaining high motion quality. The simplicity and effectiveness of RL Binding enhance the model’s ability to optimize across both modalities, resulting in improved precision and interpretability in motion generation.
$\bullet$ Comprehensive Experimental Validation: We conduct extensive experiments across multiple benchmark datasets, demonstrating the effectiveness of Motion-R1. On HumanML3D, it achieves a 3.5% improvement in MM-Dist, with Diversity and R-Precision metrics reaching state-of-the-art or on-par performance. On KIT-ML, Motion-R1 sets new records in R-Precision and FID, outperforming existing methods. On BABEL, it achieves state-of-the-art performance across all key metrics, including R-Precision, FID, MM-Dist, and Diversity. These results highlight the model’s superior ability to generate diverse, high-fidelity, and semantically aligned motions.
@@ 2 Related Work
Multimodal Large Language Models
@@
Recently, multimodal large language models (MLLMs) Wu et al. (2023); Koh et al. (2023); Tang et al. (2025) have shown impressive performance in various tasks, including text generation, reasoning, and even multimodal tasks. For example, models like GPT-4 Achiam et al. (2023) and BLIP-2 Li et al. (2023) have demonstrated the ability to understand and generate text and images. While recent MLLMs have significantly improved their perception and generation capabilities across diverse modalities, many complex real-world tasks demand not only understanding but also reasoning Banerjee et al. (2021); Xiong et al. (2024). Consequently, enhancing the reasoning abilities of MLLMs has become an important research direction.
Building on the success of CoT  Wei et al. (2022) prompting in NLP, researchers have explored extending CoT-style reasoning to multimodal settings. For instance, models like LLaVA-CoT Xu et al. (2025) introduce a novel vision-language model capable of performing autonomous, structured, and multistage reasoning by decomposing complex questions into four stages: summary, caption, reasoning, and conclusion.
While structured reasoning enhances interpretability and controllability, recent studies Tan et al. (2025); Pan et al. (2025); Liu et al. (2025) show that it can be further optimized through reinforcement learning. Building on GRPO Guo et al. (2025), we introduce a decomposed CoT paradigm for structured motion reasoning and an RL Binding mechanism, together enabling interpretable and semantically aligned motion synthesis.
Text-guided 3D Human Motion Generation
@@
Text-guided 3D human motion generation has become a key research focus in recent years. Early methods Ahn et al. (2017); Ghosh et al. (2023) primarily relied on generative adversarial networks Goodfellow et al. (2014) to synthesize human motion from text descriptions. These methods often struggled with issues such as limited diversity.
@@
To address these challenges, diffusion models Ho et al. (2020) have gained popularity in this field. Notable works Zhou & Wang (2022); Kim et al. (2023); Dabral et al. (2023) include MDM (Tevet et al., 2023), which first applies diffusion modeling to motion synthesis; MotionDiffuse (Zhang et al., 2022a) improves temporal consistency; MotionChain (Jiang et al., 2024) explores fine-grained structure and staged modeling; and MoMask (Guo et al., 2024) employs masking strategies for improved training efficiency. However, these methods often rely on complex architectures and extensive training data and are limited by their motion initialization.
@@
In parallel, VQ-VAE-based methods Guo et al. (2022a); Hong et al. (2022); Petrovich et al. (2022); Guo et al. (2022b); Athanasiou et al. (2022); Zhang et al. (2024b; 2025b); Li et al. (2025); Zhang et al. (2024d; a) discretize motion representations to facilitate efficient sequence modeling and better alignment with language, providing a more interpretable latent space. Building on this, LLMs have been integrated into the generation process to enhance the quality and control of generated motions. For instance, T2M-GPT (Zhang et al., 2023a) combines VQ-VAE with GPT in a two-stage framework for high-quality generation. Subsequent works such as MotionDiffuse Zhang et al. (2022b), MotionCLIP (Tevet et al., 2022), and MotionGPT (Jiang et al., 2023) integrate diffusion modeling, contrastive learning, and Transformer architectures to enhance fidelity and semantic consistency further. MotionAgent (Wu et al., 2024) further incorporates semantic planning for robust generalization in complex task scenarios. Meanwhile, with the development of reinforcement learning, MotionRL Liu et al. (2024) leverages PPO Schulman et al. (2017) to improve the generation based on human preferences prior knowledge, while MotionCritic Haoru Wang et al. (2025) introduces a critic model to align generated motions with textual semantics through preference-driven optimization.
Despite recent advances, T2M methods largely rely on end-to-end mapping and lack explicit reasoning over complex instructions Zhang et al. (2024c; 2025a). Our approach addresses this by combining a decomposed CoT mechanism for interpretable motion reasoning with an RL Binding strategy for efficient multimodal alignment, enabling coherent and semantically grounded motion generation without costly annotations.
@@ 3 Method
Figure 2: Overview of the Motion-R1 framework. Our method introduces two key innovations: (1) a Decomposed CoT Data Engine that generates structured motion planning traces (including <think>, <output>, and <Motion> tokens) via LLM reasoning, enabling fine-grained temporal and causal decomposition; (2) an RL Binding mechanism with GRPO-based training that streamlines optimization via embedded multi-modal alignment, ensuring semantic accuracy and motion realism without human annotations.
@@ 3.1 Motion-R1 Framework
Motion-R1 comprises two core components: a pre-trained motion tokenizer and an LLM equipped with action-oriented reasoning capabilities. The motion tokenizer discretizes continuous motion sequences into motion tokens and reconstructs them back into smooth, coherent trajectories. Meanwhile, the LLM is designed to perform structured reasoning over natural language instructions, enabling it to decompose complex action descriptions into finer-grained and logically ordered sub-actions, a process we term decomposed CoT reasoning. Based on this structured interpretation, the model generates high-quality motion token sequences that faithfully reflect the intended behavior.
As shown in Figure 2, our framework consists of two training stages. we employ the Decomposed CoT Data Engine, a novel automated pipeline for synthesizing high-quality, step-by-step reasoning data, which enables cold-start supervised tuning of the LLM to produce reasoning-augmented outputs in the <think>, <output>, and <Motion> format. By distilling the reasoning capabilities of LLMs into structured motion planning paths, this stage provides interpretable supervision and reduces reliance on costly manual annotation, enabling efficient and scalable pretraining of LLMs.
In the second stage, we refine the LLM via reinforcement learning using our proposed RL Binding strategy. Building on the GRPO Shao et al. (2024) framework, RL Binding streamlines optimization by embedding multi-modal alignment directly into the reward function, jointly evaluating embeded motion similarity with ground-truth trajectories and semantic consistency with textual descriptions. Unlike prior RL-based methods that rely on costly human annotations for preference models, RL Binding efficiently guides the model to generate motions that are both semantically faithful and motionally coherent. This design maintains simplicity and adaptability while enhancing precision, interpretability, and overall motion quality.
By combining the Decomposed CoT Data Engine and RL Binding, Motion-R1 effectively captures temporal and causal structure while ensuring semantic fidelity and motion realism, providing a unified framework for coherent, interpretable, and high-quality motion generation without the need for costly human annotations.
@@ 3.2 Motion Tokenizer
To integrate motion data, which differs significantly from natural language in structure and modality, into the LLM framework, we adopt a VQ-VAE architecture as our motion tokenizer. This approach has been widely adopted in  Zhang et al. (2023a); Jiang et al. (2023); Guo et al. (2024); Wu et al. (2024) and proven effective for 3D human motion data modeling.
The motion tokenizer comprises an encoder $E$ and a decoder $D$. Given an input motion sequence $\mathbf{m}_{1:T}\in\mathbb{R}^{T\times D}$, with $T$ frames of $D$ dimensions each, the encoder $E$ maps the sequence to a latent representation $\mathbf{z}_{1:(T/l)}\in\mathbb{R}^{(T/l)\times d}$, where $l$ is the temporal downsampling rate and $d$ is the number of latent dimensions. Each latent vector $\mathbf{z_{i}}$ is then quantized using a learnable codebook $\mathbf{C}=\{\mathbf{c}_{n}\}_{n=1}^{N}$, where $N$ is the codebook size and $\mathbf{c}_{n}\in\mathbb{R}^{d}$ represents a discrete motion token. The quantization selects the nearest code vector to each embedding:
@@
The original motion sequence is reconstructed as $\hat{\mathbf{m}}_{1:T}=D(\hat{\mathbf{z}}_{1:(T/l)})$. Following Zhang et al. (2023a), we train the VQ-VAE model using a composite objective that includes a reconstruction loss $L_{\text{reconstruct}}$, a codebook commitment loss $L_{\text{commit}}$, and an embedding loss $L_{\text{embed}}$:
@@
Here, $L_{\text{reconstruct}}$ includes a smoothed L1 loss with velocity regularization to improve generation quality. To ensure stable and efficient training, we follow  Zhang et al. (2023a) to incorporate exponential moving average (EMA) updates for the codebook along with codebook reset strategies.
3.3 Decomposed CoT Data Engine
We introduce the Decomposed CoT Data Engine, an automated module for generating structured reasoning traces to guide text-to-motion generation. This engine leverages the reasoning and semantic planning capabilities of LLMs to transform free-form motion descriptions into high-quality, step-by-step CoT action plans, providing intermediate supervision that bridges language and motion.
The engine starts by augmenting existing motion-language datasets using prompt-based LLM queries. We design comprehensive prompts that instruct the LLM to decompose tasks into logically ordered sub-actions while respecting temporal dependencies and action semantics. These prompts include clear instructions, output format constraints, and in-context examples to ensure structured reasoning outputs. The complete details of the prompt are provided in the Appendix A.3 material.
Generated CoT traces are then evaluated through an automated quality control pipeline. We employ DeepSeek-R1 Guo et al. (2025) to assess each trace for relevance, logical consistency, and conciseness. Traces that exhibit redundancy, overthinking, or verbosity are filtered and regenerated until they meet quality standards. Figure 2 illustrates the full Motion-R1 pipeline, with the Decomposed CoT Data Engine highlighted to show the structured reasoning trace generation process. For the task “A person does Tai chi”, the engine identifies the main action, decomposes it into sub-actions such as “stand up”, “arm movements”, “weight displacement”, and “hand positioning”, and further elaborates each sub-action with motion-relevant details like movement direction and involved body parts.
Each validated CoT trace is paired with its original textual description and corresponding motion sequence, forming triplets of the form (description, decomposed CoT, motion). By distilling structured action plans, the Decomposed CoT Data Engine enables the model to capture fine-grained temporal and causal dependencies directly from language, significantly improving controllability, interpretability, and generalization in complex motion generation, while drastically reducing reliance on costly manual annotation.
Figure 3: Visualization comparisons with MotionLLM Wu et al. (2024) on in-distribution and out-of-distribution prompts.